# DESY-24-134 TIF-UNIMI-2024-17 Edinburgh 2024/9

### v1

@@ Abstract
We discuss how the two existing approximate N3LO (aN3LO) sets of parton distributions (PDFs) from the MSHT20 and NNPDF4.0 series can be combined for LHC phenomenology, both in the pure QCD case and for the QCDQED sets that include the photon PDF. Using the resulting combinations, we present predictions for the total inclusive cross-section for Higgs production in gluon fusion, vector boson fusion, and associated production at the LHC Run-3. For the gluon fusion and vector boson fusion channels, the corrections that arise when using correctly matched aN3LO PDFs with N3LO cross section calculations, compared to using NNLO PDFs, are significant, in many cases larger than the PDF uncertainty, and generally larger than the differences between the two aN3LO PDF sets entering the combination. The combined aN3LO PDF sets, MSHT20xNNPDF40_an3lo and MSHT20xNNPDF40_an3lo_qed, are made publicly available in the LHAPDF format and can be readily used for LHC phenomenology.
@@ 1 Introduction
The inclusion of N3LO QCD corrections in the determination of these PDF sets is approximate in two different respects. First, N3LO perturbative evolution is only known approximately. Indeed, splitting functions are not known exactly, so an approximation to them must be constructed based on exact knowledge of their small- and large- behavior and a finite set of Mellin moments [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Knowledge of N3LO splitting functions has progressed rapidly, with more Mellin moments being gradually published over the years. In fact, now [18, 19] the full set of Mellin moments up to is available for all elements of the splitting function matrix. Heavy quark transition matrix elements at , which are needed for massless perturbative evolution in a variable-flavor number scheme, are now also fully known, albeit only recently in complete form [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Hence N3LO perturbative evolution is known exactly for all practical purposes, as we shall discuss in somewhat more detail below.
Second, however, N3LO partonic cross-sections are known only for massless deep-inelastic scattering (DIS)[31] and for the total Drell-Yan cross-section and on-shell rapidity distribution[32, 33, 34, 35, 36], which are a subset of the processes used for PDF determination. In particular, whereas the computation of the high limit of heavy quark structure functions was recently completed, permitting the determination of heavy quark transition matrix elements, the computation of corrections to DIS coefficient functions proportional to powers of of the heavy quark mass is beyond the current state of the art, as is the computation of fiducial N3LO corrections to hadronic processes, which are needed for a meaningful comparison to hadron collider data.
@@
Specifically, the uncertainty on splitting functions with Mellin moments up to for all matrix elements but was assessed in both Refs. [16, 4] with very similar results, and shown in Ref. [4] to be smaller than the uncertainty due to the missing N4LO corrections estimated by scale variation. This assessment was extended to recently in Ref. [19]. The only other uncertainty on perturbative evolution present in Ref. [4] are the contributions, computed in Ref. [30], to transition matrix elements, which however were found in Ref. [30] to agree perfectly within the uncertainty band with the previous approximation [20] adopted in Ref. [4]. Based on this, it can be safely concluded that, as mentioned, N3LO perturbative evolution is fully known for all practical purposes. Furthermore, the splitting functions of Refs. [3, 4] agree within uncertainties, with the exception of the splitting function, that differs at approximately the two-sigma level, see also the benchmarking comparison of [40, 41]. Note, however, that as far as PDF evolution is concerned is much more important than , and the effects at small- are to a large extent anti-correlated. Updated knowledge of the splitting functions effectively removes this one source of moderate discrepancy. A preliminary analysis of the effect of including these additional Mellin moments, and hence improved splitting functions, which have been determined since the original MSHT publication was presented in [42], and observed a small increase in the gluon near , though still within the original MSHT uncertainties. While this is significantly washed out in parton luminosities, it does further improve the agreement between the two PDF sets.
@@
Estimates of the impact of the use of aN3LO PDFs in the computation of the Higgs total production cross-section in both Refs. [3, 4] suggest that using NNLO PDFs, instead of aN3LO PDFs, with the N3LO matrix element for Higgs production in gluon fusion and vector boson fusion leads to an error that is comparable to, or even larger than the PDF uncertainty on the N3LO result. Also, this error is found [4] to be rather larger than the estimate of its size [43, 44] based on the corresponding shift at one less perturbative order. This suggests that using NNLO PDFs in the N3LO computation spoils the accuracy of the N3LO result, in a way that, moreover, may not be reflected in a corresponding increase in uncertainty, if the latter is estimated in this way.
@@
Given the availability of two different aN3LO PDF sets, accurate N3LO phenomenology requires therefore a comparison of the impact of the use of either of these PDF sets, and a prescription for their combination. It is the purpose of this brief note to provide a first assessment and suggest a possible prescription, having specifically in view Higgs production in gluon fusion, but presenting results also for other Higgs production modes. We will first compare the aN3LO PDFs of Refs. [3, 4], both without and with QED corrections, to each other, then assess for each of them the impact of aN3LO corrections on parton luminosities, and finally construct a combination of these PDFs based on PDF4LHC combination methodology [45]. We will finally discuss results for the cross-section for Higgs production in gluon fusion and other channels based on individual and combined aN3LO PDF sets.
@@ 2 Comparison of aN3LO PDFs
We first compare the NNPDF4.0 and MSHT20 aN3LO PDF sets both with and without QED effects.
@@
When comparing the aN3LO PDF sets, it should of course be borne in mind that the MSHT20 and NNPDF4.0 sets already differ at NNLO due to differences in dataset and methodology. On top of the difference in basic underlying methodology, namely generalized polynomial parametrization and Hessian uncertainties for MSHT, neural network parametrization and Monte Carlo uncertainties for NNPDF, several details of the theoretical approach also differ. Specifically, there are several differences in the treatment of heavy quarks: the values of heavy quark masses are somewhat different; the charm PDF is fitted to the data in NNPDF4.0, while in MSHT20 it is determined by perturbative matching conditions; and for DIS the massive and massless heavy quark schemes are matched using the Thorne-Roberts scheme by MSHT and the FONLL scheme by NNPDF. A detailed comparison of the NNLO MSHT20 and NNPDF4.0 PDF sets was performed in Ref. [2], see in particular Fig. 21 and Fig. 60 of that reference. Several of the differences seen in the comparison of Ref. [2] are also seen when comparing MSHT20 to the previous NNPDF3.1 PDF set, and indeed many of the methodological differences (in particular those related to the treatment of heavy quarks) are the same. This is relevant because MSHT20 and NNPDF3.1 were included in a detailed benchmarking presented in Ref. [47, 48], along with the CT18 PDF set. In this work, ”reduced” PDF sets from the three groups were produced by adopting common assumptions and settings (specifically for the treatment of heavy quarks), instead of the inconsistent settings used in the original fits. The reduced sets turned out to be in quite good agreement with each other within uncertainties, thereby showing that the differences between the original sets were due to these inconsistent settings.
These more consistent reduced sets were then used to produce the PDF4LHC21 combination [47]. The combination that we will discuss here is instead based on the publicly available aN3LO PDF sets, thereby leading to a more conservative uncertainty estimate that encompasses the differences between the PDF sets entering the combination, as we shall explain below. To understand the features of this combination we therefore start with a comparison of the published MSHT20 and NNPDF4.0 aN3LO sets, along with the NNLO sets, to see how much of the differences seen at aN3LO reflect those already present at NNLO, whose origin was largely understood in Refs. [47, 48].
@@
The MSHT and NNPDF aN3LO parton luminosities are compared in Figs. 4-5, respectively without and with QED effects, while, in order to appreciate better the impact of N3LO corrections, the aN3LO luminosities are compared for each set to the respective NNLO luminosities in Fig. 6, normalized to the aN3LO result. The pattern of differences reproduces that seen at the level of PDFs, with a suppression of the MSHT gluon-gluon, quark-antiquark, and gluon-quark luminosities in comparison to NNPDF in the GeV region, while the quark-quark luminosity in the two sets is very similar. The suppression can be traced to the behavior of the light quark and gluon PDFs seen in Figs. 1-2 (right). Interestingly, the difference in gluon-gluon luminosity is slightly reduced for the QED PDF sets compared to the pure QCD sets.
@@ 3 Combination of aN3LO PDFs
An effective way of arriving at a conservative uncertainty estimate in our case is to produce unweighted Monte Carlo combined PDF sets [45, 51]. This is done by first turning all the PDF sets to be combined into a set of Monte Carlo replicas, which for Hessian sets can be done using the methodology of Ref. [52], and then merging an equal number of replicas from each set into a single replica set. The merged set then corresponds to a probability distribution that is the equally likely combination of the probability distribution of the input sets, and thus will lead to uncertainties that encompass those of the sets that are being combined.
@@ 4 Implications for Higgs production
We now present predictions for Higgs production at the LHC in gluon fusion, in vector boson fusion (VBF), and in association with a weak gauge boson (). Results are collected in Tables 1-2 and displayed in Fig. 11. All results shown are computed for the LHC with TeV using N3LO matrix elements, and only differ in the input PDF set. No QED or electroweak corrections are included in the matrix element, and in particular photon–initiated processes are not included, although for e.g. associated production these enter at the percent level, see e.g. [53].
The cross-sections are computed using the ggHiggs code [44] for gluon fusion, n3loxs [43] for associated production, and proVBFH [54, 55] for vector boson fusion. The central factorization and renormalization scales are set to for Higgs production in gluon fusion, (the vector boson virtuality) for VBF, and , the event-by-event invariant mass of the system, in associated production. In Fig. 11 we do not display the cross-sections since they have the same qualitative behavior as the cross-sections, as can also be observed from Table 1.
@@
The values of of obtained for each process are also included in Tables 1-2, both for NNPDF4.0, MSHT20, and the combined PDF sets constructed here, using in each case PDFs at the required perturbative orders from the same set. For the computation of for the combined set we use as a value for the average of the cross-sections computed using NNPDF4.0 and MSHT20 NLO PDFs.
@@
Finally, we observe that for Higgs production in gluon fusion the effect of QED corrections, again as already seen in luminosities, is also not negligible, though smaller than that of using properly matched PDFs.
It is finally interesting to compare results obtained using the MC combination of aN3LO PDF sets to those that would be found by performing the textbook weighted average and uncertainty. For gluon fusion the weighted combination (with weights ) gives
@@ 5 Conclusion
In this note we have shown that the use of aN3LO PDFs is mandatory for accurate N3LO phenomenology at the LHC, focusing on the case of Higgs production cross-sections. For Higgs production in gluon fusion and in vector boson fusion the impact of switching from NNLO to aN3LO PDFs is rather larger than the PDF uncertainty, and also larger (for VBF much larger) than the difference between different PDF sets. Indeed, we have shown that differences between existing aN3LO PDF sets are moderate, and generally smaller than the difference between NNLO and aN3LO.

### v2

@@ Abstract
We discuss how the two existing approximate N3LO (aN3LO) sets of parton distributions (PDFs) from the MSHT20 and NNPDF4.0 series can be combined for LHC phenomenology, both in the pure QCD case and for the QCDQED sets that include the photon PDF. Using the resulting combinations, we present predictions for the total inclusive cross-section for Higgs production in gluon fusion, vector boson fusion, and associated production at the LHC with TeV (Run 3). For the gluon fusion and vector boson fusion channels, the corrections that arise when using correctly matched aN3LO PDFs with N3LO cross section calculations, compared to using NNLO PDFs, are significant, in many cases larger than the PDF uncertainty, and generally larger than the differences between the two aN3LO PDF sets entering the combination. The combined aN3LO PDF sets, MSHT20xNNPDF40_an3lo and MSHT20xNNPDF40_an3lo_qed, are made publicly available in the LHAPDF format and can be readily used for LHC phenomenology.
@@ 1 Introduction
The inclusion of N3LO QCD corrections in the determination of these PDF sets is approximate in two different respects. First, N3LO perturbative evolution is only known approximately. Indeed, splitting functions are not known exactly, so an approximation to them must be constructed based on exact knowledge of the small- and large- behavior at higher perturbative orders (e.g. from resummation) and a finite set of Mellin moments [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Knowledge of N3LO splitting functions has progressed rapidly, with more Mellin moments being gradually published over the years. In fact, now [18, 19] the full set of Mellin moments up to is available for all elements of the splitting function matrix. Heavy quark transition matrix elements at , which are needed for massless perturbative evolution in a variable-flavor number scheme, are now also fully known, albeit only recently in complete form [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Hence N3LO perturbative evolution is known exactly for all practical purposes, as we shall discuss in somewhat more detail below.
Second, however, N3LO partonic cross-sections are only known for a subset of the processes used for PDF determination, namely massless deep-inelastic scattering (DIS)[31, 32] and the total Drell-Yan cross-section and on-shell rapidity distribution[33, 34, 35, 36, 37]. In particular, whereas the computation of the high limit of heavy quark structure functions was recently completed, permitting the determination of heavy quark transition matrix elements, the computation of corrections to DIS coefficient functions proportional to powers of of the heavy quark mass is beyond the current state of the art, as is the computation of fiducial N3LO corrections to hadronic processes, which are needed for a meaningful comparison to hadron collider data.
@@
Specifically, the uncertainty on splitting functions with Mellin moments up to for all matrix elements but was assessed in both Refs. [16, 4] with very similar results, and shown in Ref. [4] to be smaller than the uncertainty due to the missing N4LO corrections estimated by scale variation. This assessment was extended to recently in Ref. [19]. The only other uncertainty on perturbative evolution present in Ref. [4] are the contributions, computed in Ref. [30], to transition matrix elements, which however were found in Ref. [30] to agree perfectly within the uncertainty band with the previous approximation [20] adopted in Ref. [4]. Based on this, it can be safely concluded that, as mentioned, N3LO perturbative evolution is fully known for all practical purposes. Furthermore, the splitting functions of Refs. [3, 4] agree within uncertainties, with the exception of the splitting function, that differs at approximately the two-sigma level, see also the benchmarking comparison of [41, 42]. Note, however, that as far as PDF evolution is concerned is much more important than , and the effects at small- are to a large extent anti-correlated. Updated knowledge of the splitting functions effectively removes this one source of moderate discrepancy.
A preliminary analysis of the effect of including these additional Mellin moments, and hence improved splitting functions, which have been determined since the original MSHT publication was presented in [43], and observed a small increase in the gluon near , though still within the original MSHT uncertainties. While this is significantly washed out in parton luminosities, it does further improve the agreement between the two PDF sets. Furthermore, updated PDF sets obtained using a parametrization of the N3LO splitting function constructed in [19] that includes the full information on Mellin moments (and most recently by MSHT, also the complete heavy quark transition matrix elements already used by NNPDF) now available have been presented in preliminary form both by MSHT [44, 45] and NNPDF [46]. These preliminary results allow for a comparison in which all the currently available information on perturbative evolution is consistenly used by the two groups. Based on these results, it can be concluded that, whereas the use of a common parametrization of splitting functions and transition matrix elements reduces somewhat the differences between the MSHT and NNPDF aN3LO PDF sets, results change by a small amount. We shall come back to this point in Sect. 4 below.
@@
Estimates of the impact of the use of aN3LO PDFs in the computation of the Higgs total production cross-section in both Refs. [3, 4] suggest that using NNLO PDFs, instead of aN3LO PDFs, with the N3LO matrix element for Higgs production in gluon fusion and vector boson fusion leads to an error that is comparable to, or even larger than the PDF uncertainty on the N3LO result. Also, this error is found [4] to be rather larger than the estimate of its size based on the corresponding shift at one less perturbative order [47, 48]. This suggests that using NNLO PDFs in the N3LO computation spoils the accuracy of the N3LO result, in a way that, moreover, may not be reflected in a corresponding increase in uncertainty, if the latter is estimated in this way.
@@
Given the availability of two different aN3LO PDF sets, accurate N3LO phenomenology requires therefore a comparison of the impact of the use of either of these PDF sets, and a prescription for their combination. It is the purpose of this brief note to provide a first assessment and suggest a possible prescription, having specifically in view Higgs production in gluon fusion, but presenting results also for other Higgs production modes. We will first compare the aN3LO PDFs of Refs. [3, 4], both without and with QED corrections, to each other, then assess for each of them the impact of aN3LO corrections on parton luminosities, and finally construct a combination of these PDFs based on the PDF4LHC combination methodology [49]. We will finally discuss results for the cross-section for Higgs production in gluon fusion and other channels based on individual and combined aN3LO PDF sets.
@@ 2 Comparison of aN3LO PDFs
We will compare NNPDF4.0 and MSHT PDF sets both with and without QED effects. The QED sets include a photon PDF that mixes with the quark and gluon PDFs upon combined QEDQCD evolution, and mostly differ from the pure QCD set due to slight depletion of the gluon PDF due to the momentum fraction carried by the photon. They are consequently important when assessing the behavior of the gluon PDF.
@@
When comparing the aN3LO PDF sets, it should of course be borne in mind that the MSHT20 and NNPDF4.0 sets already differ at NNLO due to differences in dataset and methodology. On top of the difference in basic underlying methodology, namely generalized polynomial parametrization and Hessian uncertainties for MSHT, neural network parametrization and Monte Carlo uncertainties for NNPDF, several details of the theoretical approach also differ. Specifically, there are several differences in the treatment of heavy quarks: the values of heavy quark masses are somewhat different; the charm PDF is fitted to the data in NNPDF4.0, while in MSHT20 it is determined by perturbative matching conditions; and for DIS the massive and massless heavy quark schemes are matched using the Thorne-Roberts scheme by MSHT and the FONLL scheme by NNPDF. A detailed comparison of the NNLO MSHT20 and NNPDF4.0 PDF sets was performed in Ref. [2], see in particular Fig. 21 and Fig. 60 of that reference. Several of the differences seen in the comparison of Ref. [2] are also seen when comparing MSHT20 to the previous NNPDF3.1 PDF set, and indeed many of the methodological differences (in particular those related to the treatment of heavy quarks) are the same. This is relevant because MSHT20 and NNPDF3.1 were included in a detailed benchmarking presented in Ref. [51, 52], along with the CT18 PDF set. In this work, dedicated PDF sets from the three groups were produced by adopting common assumptions and settings (specifically for the treatment of heavy quarks), instead of the inconsistent settings used in the original fits, as well as a relatively similar global dataset. These PDF sets were called “reduced” in Ref. [51] (though not based on a reduced dataset). These dedicated sets turned out to be in quite good agreement with each other within uncertainties, thereby showing that the differences between the original sets were partly due to these inconsistent settings.
These more consistent sets were then used to produce the PDF4LHC21 combination [51]. The combination that we will discuss here is instead based on the publicly available aN3LO PDF sets, thereby leading to a more conservative uncertainty estimate that encompasses the differences between the PDF sets entering the combination, as we shall explain below. To understand the features of this combination we therefore start with a comparison of the published MSHT20 and NNPDF4.0 aN3LO sets, along with the NNLO sets, to see how much of the differences seen at aN3LO reflect those already present at NNLO, whose origin was largely understood in Refs. [51, 52].
@@
The MSHT and NNPDF aN3LO parton luminosities are compared in Figs. 4-5, respectively without and with QED effects, while, in order to appreciate better the impact of N3LO corrections, the aN3LO luminosities (without QED) are compared for each set to the respective NNLO luminosities in Fig. 6, normalized to the aN3LO result. The pattern of differences reproduces that seen at the level of PDFs, with a suppression of the MSHT gluon-gluon, quark-antiquark, and gluon-quark luminosities in comparison to NNPDF in the GeV region, while the quark-quark luminosity in the two sets is very similar. The suppression can be traced to the behavior of the light quark and gluon PDFs seen in Figs. 1-2 (right). Interestingly, the difference in gluon-gluon luminosity is slightly reduced for the QED PDF sets compared to the pure QCD sets.
@@ 3 Combination of aN3LO PDFs
An effective way of arriving at a conservative uncertainty estimate in our case is to produce unweighted Monte Carlo combined PDF sets [49, 55]. This is done by first turning all the PDF sets to be combined into a set of Monte Carlo replicas, which for Hessian sets can be done using the methodology of Ref. [56], and then merging an equal number of replicas from each set into a single replica set. The merged set then corresponds to a probability distribution that is the equally likely combination of the probability distribution of the input sets, and thus will lead to uncertainties that encompass those of the sets that are being combined. This way if combining uncertainties leads to results that are more conservative than those that would be obtained by performing a weighted average, or possibly including correlations from the underlying data; however, correlations cannot [57] be reliably taken into account, and differences due to different but equally plausible underlying methodological assumptions are thereby correctly accounted for.
@@ 4 Implications for Higgs production
We now present predictions for Higgs production at the LHC in gluon fusion, in vector boson fusion (VBF), and in association with a weak gauge boson (). Results are collected in Tables 1-2 and displayed in Fig. 11. All results shown are computed for the LHC with TeV using N3LO matrix elements, and only differ in the input PDF set. No QED or electroweak corrections are included in the matrix element, and in particular photon–initiated processes are not included. The impact of these on the processes considered here is assessed e.g. in Ref. [58], it is negligible for gluon fusion, of order of a percent for VBF and a few percent for associate production, and it is similar for all PDF sets considered here as the photon PDF in the NNPDF4.0 and MSHT20 sets is quite similar, so it does not affect the comparisons presented here.
The cross-sections are computed using the ggHiggs code [59, 60] for gluon fusion, n3loxs [47] for associated production, and proVBFH [61, 62] for VBF. The central factorization and renormalization scales are set to for Higgs production in gluon fusion, (the vector boson virtuality) for VBF, and , the event-by-event invariant mass of the system, in associated production. In Fig. 11 we do not display the cross-sections since they have the same qualitative behavior as the cross-sections, as can also be observed from Table 1.
@@
The values of and obtained for each process are also included in Tables 1-2, both for NNPDF4.0, MSHT20, and the combined PDF sets constructed here, using in each case PDFs at the required perturbative orders from the same set. For the computation of for the combined set we use as a value for the average of the cross-sections computed using NNPDF4.0 and MSHT20 NLO PDFs.
@@
Also, we observe that for Higgs production in gluon fusion the effect of QED corrections, again as already seen in luminosities, is also not negligible, though smaller than that of using properly matched PDFs.
It is also interesting to compare results obtained using the MC combination of aN3LO PDF sets to those that would be found by performing the textbook weighted average and uncertainty. For gluon fusion the weighted combination (with weights ) gives
@@
Finally, as mentioned in the introduction, we have assessed the effect of using the preliminary variants of the aN3LO PDF sets presented here, and recently constructed by MSHT [44, 45] and NNPDF [46] using the parametrization of aN3LO splitting functions of Ref. [19], that includes all the perturbative information that is currently available, and also using for both sets the same N3LO heavy quark transition matrix element as used by NNPDF4.0. These preliminary variants have been also used in order to produce a variant combined set. Very moderate differences are found: specifically, for NNPDF, the central value of the gluon-fusion cross-section in the aN3LO baseline fit of 45.49 pb (see Table 1) becomes 45.25 pb, namely a shift of , while for MSHT the aN3LO baseline result of 44.08 pb becomes 44.26 pb, a shift of (when only updated the splitting functions the respective value is 44.38 pb, i.e. a shift of ). In the case of the VBF cross-sections, we have verified that the impact of the updated splitting functions is negligible, i.e. of order . Therefore, the difference between the central predictions of the aN3LO fits of NNPDF and MSHT is reduced by around when the same parametrization of splitting functions from [19] is used. However, the average value obtained from a combined set of PDFs is almost unchanged. Moreover, because the updated PDFs get closer, the uncertainty found with this updated combination is reduced, consistent with the reduction in theoretical uncertainty due to better knowledge of N3LO splitting functions and transition matrix elements. In summary, the updated combination leads to an almost identical central prediction and a somewhat reduced uncertainty.
Hence, we conclude that the results using the combination presented here, based on published PDF sets, can currently be relied upon to produce a reliable set of predictions and a conservative uncertainty estimate. Similarly, the individual published aN3LO PDF sets can be taken to be reliable, as updated information results in very similar predictions well within the uncertainty, which again can be taken as conservative as improved N3LO knowledge would reduce theoretical uncertainty. On the other hand, we conclude that adopting a common set of N3LO splitting function does improve the overall consistency between the MSHT20 and NNPDF4.0 aN3LO sets, especially for the gluon-gluon luminosity. As more information on N3LO terms becomes available, it will be incorporated in future PDF sets, and PDFs based on common settings could be considered for future combinations. For the time being the existing sets may be reliably used, until they will be replaced by some newer version including updates based on improvements in N3LO knowledge but also potentially on increased data and other methodological changes.
@@ 5 Conclusion
In this note we have shown that the use of aN3LO PDFs is mandatory for accurate N3LO phenomenology at the LHC, focusing on the case of Higgs production cross-sections. For Higgs production in gluon fusion and in vector boson fusion the impact of switching from NNLO to aN3LO PDFs is rather larger than the PDF uncertainty, and also larger (for VBF much larger) than the difference between different PDF sets. Indeed, we have shown that differences between existing aN3LO PDF sets are moderate, and generally smaller than the difference between NNLO and aN3LO. We have verified that updates due to improved knowledge of N3LO splitting functions and transition matrix elements does not alter this conclusion.